Find y when x = -3 for y = 1.5x + 11
1
Expert's answer
2020-04-14T16:07:25-0400

"\begin{aligned}
 & \text{Use the equation }y\text{ }=\text{ }1.5x\text{ }+\text{ }11\text{ to find the value of }y\text{ when }x\text{ is }-3. \\ 
 & \\ 
 & \text{Simply substitute }-3\text{ for }x\text{ and then evalaute}\text{.} \\ 
 & \\ 
 & y\text{ }=\text{ }1.5\left( -3 \right)\text{ }+\text{ }11 \\ 
 & \\ 
 & y\text{ }=\text{ }-4.5\text{ }+\text{ }11 \\ 
 & \\ 
 & y=11-4.5 \\ 
 & \\ 
 & y=6.5 \\ 
 & \\ 
 & \text{Therefore, value is }y=6.5. \\ 
\end{aligned}"